Caspian Sea Drinks is considering the purchase of a new water filtration system produced by Rube Goldberg Machines. This new equipment, the RGM-7000, will allow Caspian Sea Drinks to expand production. It will cost $15.00 million fully installed and will be fully depreciated over a 20 year life, then removed for no cost. The RGM-7000 will result in additional revenues of $3.22 million per year and increased operating costs of $663,005.00 per year. Caspian Sea Drinks' marginal tax rate is 29.00%. If Caspian Sea Drinks uses a 8.00% discount rate, then the net present value of the RGM-7000 is _____.
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(15,000,000/20)=$750000/year

Hence incremental cash flow=(Additional revenue-Increased operating cost)(1-tax rate)+Tax savings on Annual depreciation

=(3,220,000-663,005)(1-0.29)+(0.29*750000)

=$2032966.45/year

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=2032966.45/1.08+2032966.45/1.08^2+2032966.45/1.08^3+2032966.45/1.08^4+2032966.45/1.08^5+2032966.45/1.08^6+2032966.45/1.08^7+2032966.45/1.08^8+2032966.45/1.08^9+2032966.45/1.08^10+2032966.45/1.08^11+2032966.45/1.08^12+2032966.45/1.08^13+2032966.45/1.08^14+2032966.45/1.08^15+2032966.45/1.08^16+2032966.45/1.08^17+2032966.45/1.08^18+2032966.45/1.08^19+2032966.45/1.08^20

=19959964.28

NPV=Present value of inflows-Present value of outflows

=19959964.28-15,000,000

=$4959964.28(Approx)